A monopoly is considering selling several units of a homogeneous product as a single package. A typical consumer’s demand for the product is Qd = 50 - 0.25P, and the marginal cost of production is $120.

a. Determine the optimal number of units to put in a package.

b. How much should the firm charge for this package?
Answer #1

Answer a) Given,

Q = 50 - 0.25P

=> 0.25P = 50 - Q

=> P = (50 - Q) / 0.25

=> P = 200 - 4Q

TR (Total Revenue) = P * Q = (200 - 4Q) * Q

=> TR = 200Q - 4Q^2

MR (Marginal Revenue) = \DeltaTR / \DeltaQ

=> MR = 200 - 8Q

MC (Marginal Cost) = $120

For monopolist the profit-maximizing condition is MR = MC. So,

200 - 8Q = 120

=> 200 - 120 = 8Q

=> 8Q = 80

=> Q = 80 / 8

=> Q = 10

Here the optimal number of units for a package is 10 units.

b) From the demand function we get,

P = 200 - 4Q

=> P = 200 - (4 * 10)

=> P = $160

Therefore, here the firm should charge $160 for the package.